THe volume of air bubble at the bottom of a lake when the pressure is 3p/4 atm is 3v/8, what will be the product of the volume of the air bubble at the surface of the lake if the pressure is 2/3 of the original value and the pressure is reduced by 3/8 of the original value at constant temperature

1
Expert's answer
2021-01-19T04:36:23-0500

pV=nRT.

3p/4 x 3v/8 = 1 x 8.314 x 298.

9pV/32= 2 477,572

pV= 8 809,14489.

2/3 x 3/8 = 6/24 = 1/4.

V/4=8 809,14489.

V= 35 236,5796.